OpenStudy (anonymous):

The capacity of a storage battery, such as those used in automobile electrical systems, is rated in ampere-hours . A 50 battery can supply a current of 50 for 1.0 , or 25 for 2.0 , and so on.

OpenStudy (anonymous):

U= 3*10^6

OpenStudy (anonymous):

What volume (in liters) of gasoline has a total heat of combustion equal to the energy obtained in part (a)? (See Section 17.6; the density of gasoline is 900 .)

OpenStudy (anonymous):

Total energy supplied by a 12 V, 60 A*h battery = U=QV U=60*3600*12 Total energy supplied by a 12 V, 60 A*h battery = U=2.592*10^6 J Time required to charge the battery fully=t =Energy(U)/power Time required to charge the battery fully=2.592*10^6/0.5*10^3 Time required to charge the battery fully=5184 second Time required to charge the battery fully=1.44 hour or 1 hour 26 minute 24 second

OpenStudy (anonymous):

thanks but that is just copy and pasted from a yahoo answer and it doesnt answer my question

OpenStudy (anonymous):

yes and sorry